OpenStudy (anonymous):

I need help solving these with different methods. A. 3x-9y=3 6x-3y=-24 I need to put this one into elimination. B. 7x-3y=20 5x+3y=16 I need this one in Substitution. C. y=1/2x-6 2x+6y=19 I need this one in Elimination. An explanation would be greatly appreciated Thanks!

OpenStudy (anonymous):

A. 3x - 9y = 3 reduces to x - 3y = 1 6x - 3y = -24 reduces to 2x - y = - 8 so now we have : x - 3y = 1-->(-2) 2x - y = -8 ------------- -2x + 6y = -2 (result of multiplying by -2) 2x - y = -8 -------------add 5y = -10 y = -2 now sub -2 in for y in either of original equations x - 3y = 1 x -3(-2) = 1 x + 6 = 1 x = 1 - 6 x = -5 check.. 2x - y = -8 2(-5) - (-2) = -8 -10 + 2 = -8 -8 = -8 (correct) x = -5 and y = -2 Do you have any questions about this one ?

OpenStudy (anonymous):

The only question is why did you reduce at the beginning of the problem?

OpenStudy (anonymous):

it is not necessary to reduce...it can still be solved without reducing. I just have a habit of reducing to make things easier to solve.

OpenStudy (anonymous):

Ok, Can you help me with the 2nd and 3rd one as well?

OpenStudy (anonymous):

7x + 3y = 20 7x = -3y + 20 x = -3/7y + 20/7 now sub -3/7y + 20/7 in for x in the other equation 5x + 3y = 16 5(-3/7y + 20/7) + 3y = 16 -- distribute through the parenthesis -15/7y + 100/7 + 3y = 16 -- multiply the entire equation by 7 -15y + 100 + 21y = 112 -- subtract 100 from both sides -15y + 21y = 112 - 100 -- combine like terms 7y = 12 y = 12/7 now sub 12/7 in for y in either equation to find x 7x + 3y = 20 7x + 3(12/7) = 20 7x + 36/7 = 20 -- multiply by 7 49x + 36 = 140 -- subtract 36 from both sides 49x = 140 - 36 49x = 104 x = 104/49 check... 7x + 3y = 20 7(104/49) + 3(12/7) = 20 728/49 + 36/7 = 20 728/49 + 252/49 = 20 980/49 = 20 20 = 20 (correct) x = 104/49 y = 12/7 any questions ?

OpenStudy (anonymous):

There are times when I multiplied the entire equation by the LCD...I did this to get rid of the fractions and turn them to whole numbers

OpenStudy (anonymous):

I would help you with the last one, but I have to go somewhere...I am sorry. I will be back on later if you need more help :) Again, I am sorry

OpenStudy (anonymous):

Alright Thank you for your help. Your a lifesaver! :D